OpenStudy (anonymous):

Two cans of paint and one brush cost $37. four cans of paint and 3 bushes cost $81. Find the cost of one can of paint. a. 7 b. 37 c. 15 d. 4

OpenStudy (nikato):

Ok, let p=paint and b=brush

OpenStudy (nikato):

So now, 2p+b=37 4p+3b=81

OpenStudy (nikato):

Subtract 2p on both sides of ur first equation b=37-2p

OpenStudy (nikato):

Now substitute this value for b^ in the second equation

OpenStudy (anonymous):

So I got 4p+3(37-2p)=81. then 4p-6p=-30. then -2p=-30 and divided by -2. So the paint brushes are $15. Am i correct?

OpenStudy (nikato):

It's actual 15 per pain can. Remember p=paint and b=brush. We had just solved for p=15

OpenStudy (anonymous):

Ohh, I get it. Thanks!!

OpenStudy (nikato):

Ok, ur welcome